OpenStudy (anonymous):

Find the equation, in standard form, of the line passing through the points (3,-4) and (5,1). 5x-2y=23 y=5/2x-23/2 5x+2y=23 2x-3y=9

OpenStudy (texaschic101):

first we need to find the slope using the slope formula... slope(m) = (y2 - y1) / (x2 - x1) (3,-4) x1 = 3, y1 = -4 (5,1) x2 = 5 and y2 = 1 now lets sub slope(m) = (1 - (-4) / (5 - 3) m = (1 + 4)/2 m = 5/2 the slope is 5/2 you still with me ? Any questions so far ?

OpenStudy (anonymous):

yes, i understand the slope. i am just having trouble taking y = 5/2x+23/2 and putting it in standard form

OpenStudy (texaschic101):

Now we will use y = mx + b formula slope(m) = 5/2 (5,1) x = 5 and y = 1 now we sub y = mx + b 1 = 5/2(5) + b 1 = 25/2 + b 1 - 25/2 = b 2/2 - 25/2 = b -23/2 = b making our equation : y = 5/2x - 23/2 now we will put it in standard form y = 5/2x - 23/2 (multiply the entire equation by the LCD, which is 2) 2y = 5x - 23 (subtract 5x from both sides) -5x + 2y = -23 (now multiply the equation by -1 to make x positive) 5x - 2y = 23 <-- standard form

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

no, thank you. I understand now. :)

OpenStudy (texaschic101):

glad I could help :)